NEED HELP WITH SIMPLE ALGEBRA 1 PROBLEM! (Has to do with quadratics)
The square of a number (x^2), decreased by 6, is equal to 18 decreased by 5 times the number.

The numbers are _______ and _________.

Answer by htmentor(1343)

In equation form, the problem is:
X^2 - 6 = 18 - 5x
x^2 + 5x - 24 = 0
This can be factored as:
(x-3)(x+8) = 0
Thus the two solutions are: x = 3 and x = -8